Jesse’s Machining is looking to buy a new machine to handle a new four-year contract for machining windmill blades. The bookkeeper has provided the following information about the project. Jesse, the owner, says he will finance the machine at the local bank, which will charge 11% interest on the four-year loan. The machine requires a $1,500 annual maintenance payment, and will have a scrap value at the end of year four of zero.

Year Outflows Inflows
2019 $49,500 $0
2020 $1,500 $15,500
2021 $1,500 $17,775
2022 $1,500 $18,835
2023 $1,500 $19,984
Total $55,500 $72,094

Which of the following is correct?

A.

All answers are correct.

B.

The payback period is less than four years

C.

The NPV is $1,173

D.

The IRR is greater than the loan rate
Answer #1

Ans : Option A , all answers are correct

a.

Payback period
Project A
Year Cash flows Cummulative cash flows
0 -49500 -49500
1 14000 -35500
2 16275 -19225
3 17335 -1890
4 18484 16594
Payback period = A + B/C
a = year before where cash flow become positive
b = cummulative amount in the year where amount become positive
c = total amount of cash flow in the year where amount become positive

Payback period = 3 + 16594 / 18484

= 3 + 0.8977

= 3.8977 years

b.

NPV
Year Present value factor =1/(1+r)^n Present value factor Net Cash flows (inflows-outflows) Present value of cash flows
1 1/(1+11%) 0.9009 14000 12613
2 1/(1+11%)^2 0.8116 16275 13209
3 1/(1+11%)^3 0.7312 17335 12675
4 1/(1+11%)^4 0.6587 18484 12176
Present value of cash inflows 50673
Less: Initial outflow 49500
NPV 1173

c. IRR is trail and error method, at IRR inflows will be equal to outflows

Find the present value of cash flows at 12 % rate.
Present value factor (1/(1+rate)^n Present value factor Cash flows Present value
1/(1.12) 0.892857143 14000 12500.00
1/(1.12)^2 0.797193878 16275 12974.33
1/(1.12)^3 0.711780248 17335 12338.71
1/(1.12)^4 0.635518078 18484 11746.92
Present value of cash flows 49560
Initial investment 49500
60

IRR is approximately 12% which is more than 11%

All options are true. therefore option A is true